5 Data 6 Sales $25,000,000 7 Net operating income $3,000,000 8 Average operating assets $10,000,000 9 Minimum required rate oNeed formula and explanation
Answer #1
Data
Sales 25000000
Net Operating Income 3000000
Average Operating Assets 10000000
Minimum required rate of return 25%
Compute the ROI
Margin 12% =B7/B6
Turnover 2.5 =B6/B8
ROI 30% =B7/B8
Compute the residual income
Average Operating Assets 10000000 =B8
Net Operating Income 3000000 =B7
Minimum required rate of return 25% =B9
Residual Income 500000 =B21-B20*B22


Margin = Net Operating Income / Sales x 100
Turnover = Sales / Net Operating Income
ROI = Net Operating Income / Average Operating Assets x 100

Residual Income = Net Operating Income - Average Operating Assets x Minimum Required rate of return

1.
ROI = 30%
ROI is Net Operating income earned for Average Operating Assets employed.

2.
Since the Average Operating assets are lowered, therefore ROI increased as net operating income is associated with lower operating assets.

Data
Sales 25000000
Net Operating Income 3000000
Average Operating Assets 8000000
Minimum required rate of return 25%
Compute the ROI
Margin 12% =B7/B6
Turnover 3.125 =B6/B8
ROI 37.50% =B7/B8
Compute the residual income
Average Operating Assets 8000000 =B8
Net Operating Income 3000000 =B7
Minimum required rate of return 25% =B9
Residual Income 1000000 =B21-B20*B22


Residual Income increased by 500000, since ROI increased

3.

Data
Sales 25000000
Net Operating Income 3000000
Average Operating Assets 15000000
Minimum required rate of return 25%
Compute the ROI
Margin 12% =B7/B6
Turnover 1.666667 =B6/B8
ROI 20.00% =B7/B8
Compute the residual income
Average Operating Assets 15000000 =B8
Net Operating Income 3000000 =B7
Minimum required rate of return 25% =B9
Residual Income -750000 =B21-B20*B22


Residual Income decreased by 1250000